OpenStudy (anonymous):

Okay so we're doing Quadratic equations and i'm trying to find the vertex. i have the AOS, but how do i find the vertex if 'x' is in the 'a' position? the problem is y=x^2+6x-6

jimthompson5910 (jim_thompson5910):

what did you get for the axis of symmetry

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

how did you get - 3?

jimthompson5910 (jim_thompson5910):

that's the x coordinate of the vertex, plug that into the equation to find y

jimthompson5910 (jim_thompson5910):

y=x^2+6x-6 y=(-3)^2+6(-3)-6 y = ??

OpenStudy (anonymous):

okay i got -12

jimthompson5910 (jim_thompson5910):

close, but no

OpenStudy (anonymous):

okay, i dont know where i went wrong

jimthompson5910 (jim_thompson5910):

y=(-3)^2+6(-3)-6 y=9+6(-3)-6 y=9-18-6 y = ???

OpenStudy (anonymous):

oh..wow okay -15?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

okay thanks

jimthompson5910 (jim_thompson5910):

np